# The Impact of Large Language Models (LLMs) on Code Review Process

### v1

@@ 3.3.1 Identifying PR Events using GPT
For each PR previously labeled as GPT-assisted, we applied the heuristic H4 (explained in Table 2) to timeline events [33] to label the events that the contributor and reviewers shared as GPT-assisted or GPT-non-assisted events. This specifies when and in what capacity GPT helps with PR discussions. Figure 2 is a visual representation of our definition of the PR phases:
@@ 4. $\blacksquare$
    at_waiting_time: the time in hours spent on multi-turn conversations and feedback shared between reviewers and the author. The first waiting time is between $at\_review_{2}$ and $at\_reviewn_{n-1}$. GPT was not use for the second waiting time phase between acceptance and merge/close.
@@
To analyze the impact of LLMs on each PR phase, we mapped a chronological sequence of events (including pulling, examining, integrating, forcing the head reference, commenting on reviews, comments in lines, commit messages, and merging) to the phases. To guarantee the accuracy of our findings, we only considered PRs with sequential timestamps. In these cases, the review phase follows the submission phase, the approval step follows the review, and the merging occurs after acceptance. Finally, to ensure human participation in the review process, we excluded PRs in which the only participants, either as contributors or reviewers, were the original author and automated bots.
To have a fair comparison between the GPT-assisted and GPT-non-assisted PRs, we quantify the similarity between them by applying the Manhattan [34, 35] on the same factors as applied in RQ1 4.1. To mitigate the presence of outliers or bias in the factor, we rescale the no_commits, PR_size, no_changed_files, project_age to a range within [0, 1], applying the max-min method [35]. Since the two well-known metrics (the Manhattan and Euclidean distance produce very similar results (Spearman correlation $\rho=0.98$ [36] when computing the distances between factors), we finally test the differences between the PRs supported by or without GPT by getting the minimum distance value from the Manhattan distance.
@@ 4.1 Does GPT-assistance during the PR process reduce the resolution time?
Motivation: Recent studies have investigated the influence and popularity of ChatGPT on code review [37], [11], [24], [38]. The potential impact of using GPT models as assistants throughout the entire PR lifecycle remains largely unexplored. This research question investigates whether relying on GPT-assisted PRs affects code review resolution time in GitHub projects by analyzing the impact of GPT models on the overall merge time of PRs.
Approach: To answer this research question, we build a Multiple Linear Regression Model (MLRM) [39] to empirically explain PR lifetime. For this purpose, we examine previous studies to identify a list of known factors that significantly influence PR resolution time [4], [16], [19], [40]. Bernardo et al. [41] defines resolution time as the time it takes for a code change to be merged into the development branch after review. This implies that the change is not fully implemented until it can be built, tested, and deployed. This is possible only after merging into the main branch. We will use the time it takes for code modifications to be merged as a metric, selecting only PRs that have been merged, considering the identified factors that affect this time [40]. Furthermore, we added a two-state indicator to the dataset, with 1 representing PR that use GPT-assisted and 0 representing PR that use GPT non-assisted. This indicator is used to better understand the impact of the GPT as a factor on PR latency.
After this step, our dataset includes only the merged PRs with a total of 1,367 PRs (85.43% of the total GPT-assisted PR population as shown in Table 3) using GPT-assisted. To compare the resolution time between the GPT-assisted PRs and GPT-non-assisted PRs, we searched for the most similar PRs among the 23,874 identified GPT-non-assisted PRs by using the Manhattan distance [42] as shown in Figure 1. To compare the statistical difference between PRs using GPT-assisted or GPT-non-assisted, as an influencing factor, we use the Mann-Whitney U test, a non-parametric test that compares two independent groups that do not require normally distributed samples[43]. To carry out this statistical test, we utilize the stats package [^2] R Core Team. 2023. R: a language and environment for statistical computing. https://www.R-project.org considering a confidence level of 95% ($\alpha=0.05$). The following formula and list of factors characterize the Linear Model (LM) performed in our study:
@@ 3. $\blacksquare$
    Common factors that influence the PR latency [4, 19, 16].
@@ 4. $\blacksquare$
To mitigate the influence of skewness and enhance the clarity of our statistical examination, we implemented a logarithmic transformation on variables that exhibit a skewed distribution [44]. To implement our linear model shown in Equation 1, we follow the approach proposed by James et al. [39]. First, we build our model using the lm function that comes standard with the distribution [^3] https://www.rdocumentation.org/packages/stats/versions/3.6.2/topics/lm. Then we evaluate the degree of confidence of our model by obtaining the marginal and conditional coefficients of determination $(R^{2})$ [45]. To determine whether the adoption of GPT-assisted has a potential impact on PR latency, we check the statistical significance and the total variance of is_gpt-assisted in the model. To identify the level of significance of the variables, we rely on the traditional confidence level of $\alpha=0.05$ [46].
@@
Figure 3: Distribution of time to merge for GPT-assisted and GPT non-assisted PRs
Answer to RQ1. The results highlight the potential of GPT assistance to significantly accelerate the pull request review process, reducing merge times by more than 60%. Integrating GPT as an assistance in the development workflows can help streamline code review, improve team efficiency, and facilitate faster software delivery.
@@ 4.2 How does GPT assistance impact various phases of the PR process?
Motivation: Previous studies have shown that factors such as technology, collaboration dynamics, and project context influence the effectiveness of code reviews [4, 5, 6, 41]. More recently, AI-driven tools, particularly LLMs, have been explored for their potential to assist with PR-related development tasks and improve review efficiency [15, 14, 9]. Despite growing interest, little is known about how developers actually use GPT during the code review process or how its usage varies across different phases of a PR lifecycle. This gap motivates our investigation of the impact of GPT as assistance on review and resolution time. By identifying the specific phases in which GPT is applied and understanding its role as an assistive (rather than an automated) tool, we aim to assess where it offers the greatest benefit in the review workflow.
Approach: To investigate the impact of using GPT as assistance throughout the PR phases, we first define the distinct phases of the PR workflow, starting from the submission of a change to merge into the project branch including the presence of the waiting time, as described in detail in the section  3.3. To assess whether the use of GPT-assisted leads to statistically significant differences in PR phases, the Mann-Whitney U test was applied [47]. This non-parametric test is well-suited for comparing two independent samples, particularly when the assumption of normality cannot be guaranteed. In this study, the GPT-assisted and GPT-non-assisted datasets are treated as independent groups, as each PR belongs to only one condition (each PR is placed into only one of these two groups), with no overlapping or paired observations. Using the Mann-Whitney U test aims to determine whether the distributions of phases differ significantly between the two groups, thus providing evidence on the potential impact of GPT-assisted.
@@
Note: There is no evidence of GPT assistance being used during the submission or waiting after acceptance phases of the PR workflow.
Answer to RQ2. Our analysis reveals that GPT-assisted PRs are associated with significantly shorter durations in specific phases. The review phase is 33% faster, and the revision before acceptance is reduced by 87% compared to GPT-non-assisted PRs. However, no significant differences are observed in the change phase.
@@ 4.3 Why do contributors and reviewers use GPT during the PR process?
Motivation: Our prior analysis revealed that the use of GPT was associated with notable variations in the duration of individual PR phases, suggesting that its influence extends beyond the overall resolution time to phase-specific developer activities. While earlier studies have examined how developers use GPT for specific code review tasks [9, 12], recent work has raised concerns about its influence on review discussions and developer responses [24, 38]. Although GPT tools show promise across various software engineering activities, little is known about their role across different phases of the PR workflow. This suggests an underexplored but potentially broad impact of GPT-based tools on the code review process. To fill the gap, we analyze the purposes for which developers use GPT within GitHub PR phases.
Approach: To explore the purposes for which developers use GPT as assistance, we performed a detailed analysis of a statistical sample of 310 randomly selected GPT-assisted PRs drawn from the dataset described in Section 3.3. Initially, two authors independently examined and assigned labels to each instance (e.g., enhancement, documentation, fixing bugs). The labeling process was conducted in two rounds, each covering about 155 GPT-assisted PRs. After each round, the authors met to refine the labels, and the final classification was based on the taxonomy proposed by Tufano et al. [9]. The analysis specifically targeted events in which developers explicitly recorded their prompts to GPT and the corresponding suggestions generated by the model. This classification enabled a structured understanding of the functional contexts in which GPT-assisted PRs are integrated into the PR workflow.
@@ 1. $\blacksquare$
    Enhancement: As shown in Table 6, this is the most common activity during the stages (60.26% At Review, 42.31% Waiting before Change and 35.90% At Change), when developers offer current code as a baseline for adaptation. Most of the suggestions made by GPT are about refactoring the code, missing error handling, and making changes to existing code for performance reasons. We can see that the rewriting of the code was for minor modifications; for example, the most common modification was renaming a variable, function or class name, changing a flag variable, declaring a variable, adding/removing a conditional or reorganization a function. In contrast, we see more occasions when GPT proposes adding an error-handling instruction (try..catch). An example of a case of enhancement can be found in the PR [48], where the developer asked ChatGPT to change a Ruby plugin: updates external-posts.rb plugin, allowing the user to specify an explicit lists of urls in config.yml that are then displayed in the blog feed as external posts. 99% of the code in this change is written by gpt-4: https://chat.openai.com/share/24432d24-36a7-4d6f-a5c0-d7e5142f68cd
@@ 2. $\blacksquare$
    Implementation: Contrary to our expectations, in our sample we observe that the Implementation in which the developers ask GPT to generate a piece of code (from scratch) has a low frequency, as shown in Table 6 where the ratio is split on 6.41% At Review, 10.26% Waiting before Change and 8.97% At Change. An illustrative instance is presented in [49], where a developer sought assistance from GPT in generating code to retrieve remote resources, and GPT responded with a Python-based implementation. We identified several cases in which reviewers expressed disagreement with the solutions generated by GPT. One such example is documented in [50], where the reviewer remarked, ”I will not blindly accept any AI-generated code” and ”using GPT so might not work as intended.”.
@@ 3. $\blacksquare$
    Bug fixing:. This task ranks as the second most common use case for GPT assistance, occurring primarily during the Review phase (25.64%) and to a lesser extent during the Change phase (21.80%). In many instances, developers consult GPT to diagnose why a specific piece of code is not functioning correctly or to request help in fixing an error (e.g., “GPT found 2 config files not exist in the dataset list.”) [51]. GPT-generated suggestions also include improvements such as adding error-handling logic —for example: “The code does not seem to have any error handling. This could lead to unexpected behavior if an error occurs. Consider adding try-catch blocks to handle potential errors.” [52].
@@ 4. $\blacksquare$
    Testing: We find few cases where developers ask for GPT-assisted about testing related tasks in our sample (every single character of tests is generated by GPT-4) [53], (I have added the shorthand option -f for the –files flag in the _ask.py file and updated the test case in test_gpt_cli.py.)[27]. We also notice that the testing inquiries to GPT were for more complex tasks such as classes and core logic testing (Chat GPT says this might work: the class AZDemoContentTest extends QuickstartFunctionalTestBase…)[54].
@@ 5. $\blacksquare$
    Documentation: We observe that developers utilize GPT both for generating new documentation — such as implementation descriptions in markdown files or tutorials on how to use specific functionalities [55] — and for modifying existing documentation, including improving navigability, enhancing readability and formatting, and updating outdated links [56]. Notably, in most instances, GPT-generated suggestions pertain to the README (.md) file [57, 58].
@@ 6. $\blacksquare$
    Other tasks: We notice that At Changed Phase and At Waiting before Change, non-programming related tasks count for the most requested activity with 28.20% and 17.95% respectively. Among the most common questions developers ask GPT for suggestions are about translation, file versions, and setting up CD/CI workflows (YAML files, file configuration, Docker, Kubernetes), Copyright, and User Interfaces [59, 60, 61].
Table 6: Frequency of Tasks That Developers Ask GPT for Assistance in PR Phases Within a Sample of 300 GPT PRs
Tasks At Review % At Waiting for change % At Change %
Enhancement 60.26 42.31 35.90
Implementation 6.41 10.26 8.97
Bug Fix 25.64 15.38 21.80
Testing 1.28 0.00 5.13
Documentation 11.54 12.82 10.26
Other 12.82 17.95 28.20

### v2

@@ 3.3.2 Identifying PR Events using GPT
For each PR previously labeled as GPT-assisted, we applied the heuristic H4 (explained in Table 1) to timeline events [GitHubTimeline] to label the events that the contributor and reviewers shared as GPT-assisted or GPT-non-assisted events. This specifies when and in what capacity GPT helps with PR discussions. Figure 2 is a visual representation of our definition of the PR phases:
@@ 4. $\blacksquare$
    at_waiting_time: refers to the duration, in hours, spent in iterative exchanges and feedback between reviewers and the contributor. The first interval (waiting before change) captures the time between $at\_review_{2}$ and $at\_review_{n-1}$, while the second (waiting after acceptance) represents the period from acceptance to merge. During the waiting phases, several iterative exchanges occur between the contributor and the reviewer.
It is important to note that when a second review (event = reviewed) is absent in the PR timeline, the PR transitions directly to the At Change phase, indicating that no multiple back-and-forth interactions occurred between the contributor and the reviewers. Furthermore, We did not identify any events that indicate any GPT-assisted activity during the Submission and Waiting after Acceptance phases; therefore, these phases were excluded from further analysis in our study.
@@
To analyze the impact of LLMs on each PR phase, we mapped a chronological sequence of events (including pulling, examining, integrating, forcing the head reference, commenting on reviews, comments in lines, commit messages, and merging) to the phases. To maintain the validity of our results, only pull requests with sequential timestamps were included in the analysis. In these cases, the review phase follows the submission phase, the approval step follows the review, and the merging occurs after acceptance. Finally, to ensure human participation in the review process, we excluded PRs in which the only participants, either as contributors or reviewers, were the original author and automated bots.
@@ 4.1 Does GPT-assistance during the PR process reduce the resolution time?
Motivation: Recent studies have investigated the influence and popularity of ChatGPT on code review [guo2024exploring], [grewal2024analyzing], [hao2024empirical], [watanabe2024use]. The potential impact of using GPT models as assistants throughout the entire PR lifecycle remains largely unexplored. This research question investigates whether relying on GPT-assisted PRs affects code review resolution time in GitHub projects by analyzing the impact of GPT models on the overall merge time of PRs.
Approach: To answer this research question, we build a Multiple Linear Regression Model (MLRM) [james2013introduction] to empirically explain PR lifetime. For this purpose, we examine previous studies to identify a list of known factors that significantly influence PR resolution time [zhang2022pull], [moreira2021factors], [rose2017towards], [silva2020measuring]. bernardo2018studying defines resolution time as the time it takes for a code change to be merged into the development branch after review. This implies that the change is not fully implemented until it can be built, tested, and deployed. This is possible only after merging into the main branch. We will use the time it takes for code modifications to be merged as a metric, selecting only PRs that have been merged, considering the identified factors that affect this time [silva2020measuring]. After this step, our dataset includes only the merged PRs with a total of 1,367 PRs (85.43% of the total GPT-assisted PR population as shown in Table 7) using GPT-assisted. To compare the resolution time between the GPT-assisted PRs and GPT-non-assisted PRs, we used the GPT-non-assisted PRs dataset explained in section 3.3.1 (Finding Similarity GPT-non-assisted PRs with GPT-Assisted PRs). Furthermore, we added a two-state indicator to the dataset, with 1 representing PR that use GPT-assisted and 0 representing PR that use GPT non-assisted. This indicator is used to better understand the impact of the GPT as a factor on PR latency.
To compare the statistical difference between PRs using GPT-assisted or GPT-non-assisted, as an influencing factor, we use the Mann-Whitney U test, a non-parametric test that compares two independent groups that do not require normally distributed samples[nachar2008mann]. To carry out this statistical test, we utilize the stats package [^2] R Core Team. 2023. R: a language and environment for statistical computing. https://www.R-project.org considering a confidence level of 95% ($\alpha=0.05$). The following formula and list of factors characterize the Linear Model (LM) performed in our study:
@@ 3. $\blacksquare$
    Common factors that influence the PR latency [zhang2022pull, rose2017towards, moreira2021factors].
@@ 4. $\blacksquare$
To mitigate the influence of skewness and enhance the clarity of our statistical examination, we implemented a logarithmic transformation on variables that exhibit a skewed distribution [feng2014log]. To implement our linear model shown in Equation 1, we follow the approach proposed by james2013introduction. First, we build our model using the lm function that comes standard with the distribution [^3] https://www.rdocumentation.org/packages/stats/versions/3.6.2/topics/lm. Then we evaluate the degree of confidence of our model by obtaining the marginal and conditional coefficients of determination $(R^{2})$ [nakagawa2017coefficient]. To determine whether the adoption of GPT-assisted has a potential impact on PR latency, we check the statistical significance and the total variance of is_gpt-assisted in the model. To identify the level of significance of the variables, we rely on the traditional confidence level of $\alpha=0.05$ [chavalarias2016evolution].
Following the methodology of  cole2017statistics, we quantified the percentage reduction between GPT-assisted and non-assisted PRs over the entire PR lifecycle (i.e., resolution or merge time), expressing the difference relative to the median merge time of non-assisted PRs:
[…]
@@
Figure 3: The distribution of merge times reveals that GPT-assisted PRs have a median merge time of approximately 9 hours, compared to 23 hours for non-assisted PRs.
Answer to RQ1. The results highlight the potential of GPT assistance to significantly accelerate the pull request review process, with percentage of reduction of the merge times by 61%. Integrating GPT as an assistance in the development workflows can help streamline code review, improve team efficiency, and facilitate faster software delivery.
@@ 4.2 How does GPT assistance impact various phases of the PR process?
Motivation: Previous studies have shown that factors such as technology, collaboration dynamics, and project context influence the effectiveness of code reviews [zhang2022pull, kudrjavets2022mining, baysal2016investigating, bernardo2018studying]. More recently, AI-driven tools, particularly LLMs, have been explored for their potential to assist with PR-related development tasks and improve review efficiency [davila2024tales, xiao2024generative, tufano2024unveiling]. Despite growing interest, little is known about how developers actually use GPT during the code review process or how its usage varies across different phases of a PR lifecycle. This gap motivates our investigation of the impact of GPT as assistance on review and resolution time. By identifying the specific phases in which GPT is applied and understanding its role as an assistive (rather than an automated) tool, we aim to assess where it offers the greatest benefit in the review workflow.
Approach: To investigate the impact of using GPT as assistance throughout the PR phases, we first define the distinct phases of the PR workflow, starting from the submission of a change to merge into the project branch including the presence of the waiting time, as described in detail in the section  3.3 (Mapping Events to PR Phases). To assess whether the use of GPT-assisted leads to statistically significant differences in PR phases, the Mann-Whitney U test was applied [mann1947test]. This non-parametric test is well-suited for comparing two independent samples, particularly when the assumption of normality cannot be guaranteed. In this study, the GPT-assisted and GPT-non-assisted datasets are treated as independent groups, as each PR belongs to only one condition (each PR is placed into only one of these two groups), with no overlapping or paired observations. Using the Mann-Whitney U test aims to determine whether the distributions of phases differ significantly between the two groups, thus providing evidence on the potential impact of GPT-assisted. We computed percentage reductions in PR phases between GPT-assisted and GPT non-assisted cases, following the definition of  cole2017statistics as defined in Formula 2. To account for the asymmetry of the raw percentage measures, we also calculated percentage of reduction in logarithmic transformation, confirming a substantial time benefit associated with GPT assistance: $109.9\%$ and $208.7\%$ reductions for the At Review and At Waiting for Change phases, respectively.
@@
Note: There is no evidence of GPT assistance being used during the submission or waiting after acceptance phases of the PR workflow. Consequently, these phases are excluded from the table.
Answer to RQ2. Our analysis reveals that GPT-assisted PRs are associated with significantly shorter durations in specific phases. The review phase is 66.7% faster, and the revision before acceptance is reduced by 87.5% compared to GPT-non-assisted PRs. However, no significant differences are observed in the change phase.
@@ 4.3 Why do contributors and reviewers use GPT during the PR process?
Motivation: Our prior analysis revealed that the use of GPT was associated with notable variations in the duration of individual PR phases, suggesting that its influence extends beyond the overall resolution time to phase-specific developer activities. While earlier studies have examined how developers use GPT for specific code review tasks [tufano2024unveiling, chouchen2024so], recent work has raised concerns about its influence on review discussions and developer responses [hao2024empirical, watanabe2024use]. Although GPT tools show promise across various software engineering activities, little is known about their role across different phases of the PR workflow. This suggests an underexplored but potentially broad impact of GPT-based tools on the code review process. To fill the gap, we analyze the purposes for which developers use GPT within GitHub PR phases.
Approach: To examine the purposes for which developers use GPT as assistance, we analyzed a statistical sample of 310 randomly selected PRs with a confidence level of 95% and a margin error of 5%, drawn from the dataset of 1,600 GPT-assisted PRs as described in Section 3.2.3. The labeling followed the taxonomy proposed by tufano2024unveiling and was conducted in two rounds of 155 PRs each. Two authors independently assigned labels (enhancement, implementation, bug fix, testing, documentation, other) to the instances, then met after each round to refine definitions. The authors resolved any discrepancies or disagreements in the labeling through discussion and consensus, rather than the formal calculation of inter-rater reliability metrics This approach allowed the authors to assign labels consistently and accurately even in nuanced cases. The annotators agreed that in some cases, multiple labels should be applied to one task, even if it is considered ”atomic”. The analysis specifically targeted events in which developers explicitly recorded their requests to GPT and the corresponding suggestions generated by the model. This classification enabled a structured understanding of the functional contexts in which GPT-assisted PRs are integrated into the PR workflow.
@@ 1. $\blacksquare$
    Enhancement: As shown in Table 4, this is the most common activity during the stages (60.26% At Review, 42.31% Waiting before Change and 35.90% At Change), when developers offer current code as a baseline for adaptation. Most of the suggestions made by GPT are about refactoring the code, missing error handling, and making changes to existing code for performance reasons. We can see that the rewriting of the code was for minor modifications; for example, the most common modification was renaming a variable, function or class name, changing a flag variable, declaring a variable, adding/removing a conditional or reorganization a function. In contrast, we see more occasions when GPT proposes adding an error-handling instruction (try..catch). An example of a case of enhancement can be found in the PR [pullrequest2059], where the developer asked ChatGPT to change a Ruby plugin: updates external-posts.rb plugin, allowing the user to specify an explicit lists of urls in config.yml that are then displayed in the blog feed as external posts. 99% of the code in this change is written by gpt-4: https://chat.openai.com/share/24432d24-36a7-4d6f-a5c0-d7e5142f68cd
@@ 2. $\blacksquare$
    Implementation: Contrary to our expectations, in our sample we observe that the Implementation in which the developers ask GPT to generate a piece of code (from scratch) has a low frequency, as shown in Table 4 where the ratio is split on 6.41% At Review, 10.26% Waiting before Change and 8.97% At Change. An illustrative instance is presented in [pullrequest1858], where a developer sought assistance from GPT in generating code to retrieve remote resources, and GPT responded with a Python-based implementation. We identified several cases in which reviewers expressed disagreement with the solutions generated by GPT. One such example is documented in [pullrequest188], where the reviewer remarked, ”I will not blindly accept any AI-generated code” and ”using GPT so might not work as intended.”.
@@ 3. $\blacksquare$
    Bug fixing:. This task ranks as the second most common use case for GPT assistance, occurring primarily during the Review phase (25.64%) and to a lesser extent during the Change phase (21.80%). In many instances, developers consult GPT to diagnose why a specific piece of code is not functioning correctly or to request help in fixing an error (e.g., “GPT found 2 config files not exist in the dataset list.”) [pullrequest1171]. GPT-generated suggestions also include improvements such as adding error-handling logic —for example: “The code does not seem to have any error handling. This could lead to unexpected behavior if an error occurs. Consider adding try-catch blocks to handle potential errors.” [pullrequest238].
@@ 4. $\blacksquare$
    Testing: We find few cases where developers ask for GPT-assisted about testing related tasks in our sample (every single character of tests is generated by GPT-4) [pullrequest4616], (I have added the shorthand option -f for the –files flag in the _ask.py file and updated the test case in test_gpt_cli.py.)[pullrequest43]. We also notice that the testing inquiries to GPT were for more complex tasks such as classes and core logic testing (Chat GPT says this might work: the class AZDemoContentTest extends QuickstartFunctionalTestBase…)[pullrequest3108].
@@ 5. $\blacksquare$
    Documentation: We observe that developers utilize GPT both for generating new documentation — such as implementation descriptions in markdown files or tutorials on how to use specific functionalities [pullrequest109] — and for modifying existing documentation, including improving navigability, enhancing readability and formatting, and updating outdated links [pullrequest28300]. Notably, in most instances, GPT-generated suggestions pertain to the README (.md) file [pullrequest36, pullrequest736].
@@ 6. $\blacksquare$
    Other tasks: We notice that At Changed Phase and At Waiting before Change, non-programming related tasks count for the most requested activity with 28.20% and 17.95% respectively. Among the most common questions developers ask GPT for suggestions are about translation, file versions, and setting up CD/CI workflows (YAML files, file configuration, Docker, Kubernetes), Copyright, and User Interfaces [pullrequest27905, pullrequest1122, pullrequest31].
Table 4: Frequency of Tasks That Developers Ask GPT for Assistance in PR Phases Within a Sample of 310 GPT PRs
Tasks At Review % At Waiting for change % At Change % Enhancement 60.26 42.31 35.90 Implementation 6.41 10.26 8.97 Bug Fix 25.64 15.38 21.80 Testing 1.28 0.00 5.13 Documentation 11.54 12.82 10.26 Other 12.82 17.95 28.20
Note: Tasks may overlap across phases. A single GPT-assisted PR can involve multiple task types (e.g., enhancement and bug fix) and appear in more than one phase (review, waiting for change, or change). Therefore, percentages are not mutually exclusive and may sum to more than 100%.